ganeshie8 (ganeshie8):

\(f(x) = x^2 + \pi x + 1\) and \(\alpha, \beta\) are zeroes of \(f(x)\) find the value of \((\alpha + 1)(\beta + 1)\)

ganeshie8 (ganeshie8):

@Jhannybean here is a challenging question that can be tried using vietas formula

OpenStudy (aaronandyson):

what is "vietas formula"?

OpenStudy (jhannybean):

Oh so first you would do the product rule

OpenStudy (aaronandyson):

its Aaron @Nnesha and not aaroon.

OpenStudy (anonymous):

\[ (x-\beta)(x-\alpha)=x^2-\alpha x-\beta x+\alpha\beta=x^2+(-\alpha-\beta)x+\alpha\beta \]

ganeshie8 (ganeshie8):

^^

OpenStudy (jhannybean):

Bleh, confused.

ganeshie8 (ganeshie8):

@AaronAndyson wio's reply pretty much summarizes the vieta's formula, just compare the coefficients both sides of equation in wio's reply

OpenStudy (jhannybean):

Ohh, wait wait wait

OpenStudy (aaronandyson):

@Nnesha it's okay, nest time kindly take care.

OpenStudy (anonymous):

\[x^2+(-\alpha-\beta)x+\alpha\beta \]Corresponds to \[ x^2+\pi x+1 \]

OpenStudy (jhannybean):

Dangit wio.

OpenStudy (aaronandyson):

thanks @ganeshie8

OpenStudy (jhannybean):

Yeahh that was one of the forms wasn't it.

OpenStudy (anonymous):

Then remember that: \[ (\alpha+1)(\beta+1)= \alpha\beta + \alpha+\beta+1 \]I don't even know what the formula is.

OpenStudy (anonymous):

Does it make sense now @Jhannybean ?

OpenStudy (jhannybean):

Kind of,

OpenStudy (anonymous):

So what is the answer?

OpenStudy (jhannybean):

-.-

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[x^2+(-\alpha-\beta)x+\alpha\beta \]Corresponds to \[ x^2+\pi x+1 \] \(\color{blue}{\text{End of Quote}}\) Trying to understand this atm.

ganeshie8 (ganeshie8):

compare apples to apples and coconuts to coconuts its an art

OpenStudy (anonymous):

Think of them as being dot products with the vector \(\langle x^2,x,1\rangle\).

ganeshie8 (ganeshie8):

\[x^2+\color{red}{(-\alpha-\beta)}x+\color{blue}{\alpha\beta} \] Ccprresponds to \[x^2+\color{red}{\pi} x+\color{blue}{1}\] compare the color coded parts

OpenStudy (jhannybean):

Can you elaborate a little more on how I to compare them?

OpenStudy (jhannybean):

to*

Nnesha (nnesha):

hmm that is quadratic equation use headphone method find two number if you multiply them you should get 1 and if you add or subtract them you should get \[\pi \] :P :D that's all i know

OpenStudy (anonymous):

i got this

ganeshie8 (ganeshie8):

yes @PFEH.1999 the answer should look something simpler like that.. il confirm one sec..

OpenStudy (unklerhaukus):

For a quadratic equation of the form \[y(x) =ax^2+bx+c\] With roots \(\alpha,\beta\) the sum of the roots (zeroes):\[\alpha+\beta=-b/a\] and the product of the roots:\[\alpha\beta=c/a\] For the quadratic \[f(x)=x^2+\pi x+1\]\[a=1,\quad b=\pi,\quad c=1\]\[\Downarrow\]\[\alpha+\beta = -\pi,\qquad\alpha\beta=1\] \[(\alpha+1)(\beta+1)= \alpha\beta + (\alpha+\beta)+1 \\ \qquad\qquad=1-\pi+1 \\\qquad\qquad= 2-\pi\]

OpenStudy (anonymous):

yes i used @UnkleRhaukus method fo alpha + beta and alpha.beta...i forgot that alpha + beta = -b/a i used -b/2a :D

ganeshie8 (ganeshie8):

\[\Uparrow\]

ganeshie8 (ganeshie8):

Ah i see how you got fractions ;) @Jhannybean does Unkles answer make sense more or less..

OpenStudy (jhannybean):

@UnkleRhaukus ' explanation makes things easier to understand. It's like Vieta's Formula for dummies.

ganeshie8 (ganeshie8):

you're right "vieta's formula" is just a fancy name for that method ;p

OpenStudy (jhannybean):

Sorry I'm so slow at this... still learning

ganeshie8 (ganeshie8):

Its all good :D Next problem going to be a bit more challenging !

Nnesha (nnesha):

:(

OpenStudy (jhannybean):

Yay

OpenStudy (jhannybean):

Now I understand the method of computation for these problems. That helps me understand where I should start. Lol.

OpenStudy (kainui):

Since we have the roots, we can just write it as factored since plugging in alpha or beta will give us 0.\[\Large x^2+\pi x+1=(\alpha - x) (\beta - x)\] Now all we need to do is plug in -1 to get: \[\Large (-1)^2+\pi (-1)+1=(\alpha +1) (\beta +1)\] So there we have it, -pi+2.

ganeshie8 (ganeshie8):

Clever xD

OpenStudy (kainui):

Merry christmas =P

ganeshie8 (ganeshie8):

Merry christmas :) have good time Kai !